Betting Terms
This policy applicable to our website https://tot-bet.com/. Please read it carefully because they form a legal agreement between you and us in relation to access and use of our website, together with other terms located in the respective section.
If the outcome of a game cannot be verified by credible means, we reserve the right to delay calculation of the winnings until the official confirmation. If we find out that the outcome of a game is displayed incorrectly, we reserve the right to correct it anytime and alter the winnings accordingly. In case of any discrepancies, rules governing respective sport below shall prevail over the general regulations.
We accept bets only through designated sections of our online platform. Bets are not accepted in any other form (e.g. via email, phone, fax, customer support etc.). Even if the bet was received, it will be automatically void.
You can cancel a bet only if it remains fully unmatched. Your cancellation of an unmatched bet becomes effective once we send you confirmation of the cancellation. If your cancellation request is not received and processed for whatever reason, you acknowledge that your bet may remain outstanding and available for acceptance.
We reserve the right to refuse the whole or any part of any bet request for any reason and at our sole discretion.
We reserve the right to impose minimum and maximum amount for bets and change them from time to time without prior notice.
We reserve the right to limit the winning per bet or impose the winning limit provided by the law. This limit may be updated from time to time. The maximum winning per bet is limited to :currency 1,000,000. In the event a winning is issued for more than the said winning limit, the limit shall apply and winning shall be capped at :currency 1,000,000.
We reserve the right to void any bet:
- If bets were mistakenly accepted when the outcome of game was already known.
- In the case of any obviously incorrectly displayed or calculated prices (deviation of more than 100% compared to the average potential winning).
- The sport event itself was held with a violation (including without limitations unusual period length, counting procedure, format of a match etc.).
- If bets were placed when the account did not have sufficient funds to cover the respective bet. If an account has insufficient funds as a result of a deposit that has been cancelled by the payment processing party, we reserve the right to cancel any bet that may have been accepted retroactively.
- If bets were placed with a conflict of interest.
- If bets were mistakenly accepted when the respective market already closed.
- If bets were placed on selection that was declared non-runner/injured or did not participate in the game for any other reason.
- In other cases provided in respective sport rules below.
- In other cases where there is a chance that the game was held with violations or the results could have been known or were annulled.
If one or more of the multiple bets is voided, they are discounted from the multiple selection, e.g. if it was a five-fold selection and one bet got voided, it becomes as four-fold selection.
We offer following types of betting:
"Single" is wagering on a single event. Win on "single" bet is the bet amount multiplied by the odds of the event.
"Parlay" is wagering on several events that do not correlate with each other. The winnings for "parlay" are settled as the bet amount multiplied by the product of the odds of all the events included in the parlay. You can include any combination of outcomes for any sport events that do not correlate with each other in parlays. "Parlays" are considered as winners if all events in parlays are predicted correctly. If at least one event is not predicted correctly, the parlay is lost. The maximum odds of "Parlay" bet are 2000.
"System" is a full combination of parlays of the same dimension chosen by you. "System" winnings are settled as the amount of winning parlays that are included in the "system". The maximum odds of "System" bet is 2000.
"Parlay+" is a daily additional list of events where bets are accepted with increased odds on Special terms.
Betting through "Parlay+" allows increasing "Parlay" odds compared to the usual line. It also increases your chances while he is playing with the maximum "Parlay" odds (K = 2000) by reducing the number of events.
- The maximum bet amount is equivalent of $500.
- The total number of events in one "Parlay" should be not less than three.
- The maximum odds of "parlay" are 2000.
- "System" is not accepted.
In "Parlay" and "System" betting, it is not allowed to include different outcomes of the same event (match winner, tournament winner) even if they are not directly connected. If incorrect forms of betting are included in parlays or systems, the bets will be refunded even if the computer software has accepted the wager.
Final odds for winnings on all listed betting types are determined taking into account the specifics of determining odds for certain types of outcomes (hitting the spread on the handicap or total, etc.), postponement of the event or the transfer of the event to another date. The transfer time is defined in betting rules. In these events, the odds for winnings are "1".
- Bets are accepted on home wins (marked as "1" in lines), draw (marked as "X" in lines) and away wins (marked as "2" in lines). The bet is a winner if you have predicted the outcome correctly.
- Home win or draw - marked "1X". The bet is a winner if the home team wins or the game ends in a draw. No draw (Home or away win) - marked "12". The bet is a winner if the game does not end in a draw. Away win or draw - marked "X2". The bet is a winner if the away team wins or the game ends in a draw.
- You can bet on the matches with handicaps (marked H). After the match, the defined handicap adds to the score of the chosen team. If the score for the chosen team is correctly predicted after the addition of the handicap, the bet is a winner. If the score is drawn after the handicap has been added, the bet will be refunded. In parlays, it will count as odds "1". If the other team wins after the addition of the handicap, the bet is a loser.
- Total - the sum of goals scored during the match. To make a winning bet, you must correctly predict over (Ov) or under (Un) X goals that will be scored during the match. Individual total (iTotal) - how many goals the chosen team will score. You must bet on Over or Under. If the score is exactly on spread, the bet will be refunded, and parlays will be settled at the odds "1".
- Qualifying to the next round - bet on the team to be qualified to the next round of contests (tournament, cup, etc.) is available.
- Results of the participants. In such betting, you predict whether a participant reaches a certain stage of the competition (e.g., 1/8, 1/4, 1/2, finals, etc.), or what place the participant will take in the tournament (group, etc.). If the declared participant in the tournament for some reason did not participate in the competition, all bets on the event with his participation are settled at the odds "1".
- In football, basketball and ice hockey matches, various handicap and total values are offered for betting.
- In the game "who is higher", it is proposed to determine which team will take higher place in the result table of the competition. Additional terms and conditions on this type of bets may be specified in the event information.
-
Home - Away.
- Home or away team win in the selected matches with handicaps. Win is determined by the difference of goals (points) scored by the home and away teams taking into account handicap.
- Predict whether the number of scored goals (points) by home team and away team in the selected matches will be more or less than the proposed number.
- Number of home team or away team wins, draws in the selected matches. There are other types of bets. If one or more of the proposed matches did not take place, was interrupted or not finished, the bets on "Home - Away" are settled at the odds "1".
- Double Asian Handicap is a bet on the game with a handicap in which the Handicap value (H) is multiple of 0.25, but not multiple of 0.5, for example: H = -0.25, +0.25, -0.75, +0.75, etc. Such a bet is interpreted as two (simple, "half") bets with the same odds and with the nearest common Handicap value (H1 = H - 0.25 and H2 = H + 0.25). The amount of each "half" bet is equal to the half of the "Double" bet amount.
-
Double Total Bet is a Total bet in which the Total value (T) is multiple of 0.25, but
not multiple of 0.5, for example: T = 2.25, 2.75, 3.25, etc. Such a bet is interpreted
as two (simple, "half") bets at the same odds and with the nearest common Total value
(T1 = T - 0.25 and T2 = T + 0.25). The amount of each "half" bet is equal to the half
of the "Double" bet amount.
The following examples are simplified four options of settlement for such "Double" bets:
Option 1. Both half bets are won by the user. In this case, the payout (S) is S=k*C
Where k is the odds listed in line
C is the sum of the user's bet.
To settle the "Parlay", the odds K = k is used
Option 2. One half of the bet is won and the other is refunded. In this case, S=(k+1)/2*C, and to settle the "Parlay", the odds K=(k +1)/2 will apply Option 3. One half of the bet is refunded, and the other is lost. In this case, S=0.5xC, and to settle the "Parlay bets", the odds K=0.5 will apply Option 4. Both bets are lost. In this case, S=0, and the odds to settle the "Parlay" is K=0
Example 1 Team Double Asian Handicap Odds Bavaria: -0.25 (0, -0.5) 2.0 Juventus: +0.25 (0, +0.5) 1.8 If the "double Asian handicap" bet of $ 200 was placed on: Bavaria Madrid
Bavaria win: both "halves" of the initial "double Asian handicap" bet win (0 and -0.5).
Payout: (100 * 2.0 + 100 * 2.0) = $ 400, and the bet odds in the "Parlay" are equal to 400/200 = 2.0.Bavaria draw: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 0.0 + 100 * 1.0) = $ 100, and the bet odds in the "Parlay" are equal to 100/200 = 0.5.Bavaria lose: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.If the "double Asian handicap" bet of $ 200 was placed on: Juventus
Juventus win: both halves of the initial "double Asian handicap" bet win (0 and +0.5).
Payout: (100 * 1.8 + 100 * 1.8) = $ 360, and the bet odds in the "Parlay" are equal to 360/200 = 1.8.Juventus draw: you lose one "half" of the initial "double Asian handicap" bet and other "half" bet has the result of a "return".
Payout: (100 * 1.8 + 100 * 1.0) = $ 280, and the bet odds in the "Parlay" are equal to 280/200 = 1.4.Juventus lose: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.Example 2 Team Double Asian Handicap Odds Bavaria: -0.25 (0, -0.5) 2.0 Juventus: +0.25 (0, +0.5) 1.8 If the "double Asian handicap" bet of $ 200 was placed on: Bavaria Madrid
Bavaria win with a difference of 2 goals or more: both "halves" of the initial "double Asian handicap" bet win (-0.5 and -1).
Payout: (100 * 2.0 + 100 * 2.0) = $ 400, and the bet odds in the "Parlay" are equal to 400/200 = 2.0.Bavaria win with a difference of 1 goal: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 2.0 + 100 * 1.0) = $ 300, and the bet odds in the "Parlay" are equal to 300/200 = 1.5Bavaria lose or draw: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.If the "double Asian handicap" bet of $ 200 was placed on: Juventus
Juventus wins or draw: both "halves" of an initial "double Asian handicap" bet win (+0.5, +1).
Payout: (100 * 1.8 + 100 * 1.8) = $ 360, and the bet odds in the "Parlay" are equal to 360/200 = 1.8.Juventus loses with a difference of 1 goal: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 0.0 + 100 * 1.0) = $ 100, and the bet odds in the "Parlay" are equal to 100/200 = 0.5.Juventus loses with a difference of 2 goals or more: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.Example 3 Team Double Asian Handicap Odds Bavaria: -1.25 (-1, -1.5) 2.0 Juventus: +1.25 (+1, +1.5) 1.8 If the "double Asian handicap" bet of $ 200 was placed on: Bavaria Madrid
Bavaria win with a difference of 2 goals or more: both "halves" of the initial "double Asian handicap" bet win (-1 and -1.5).
Payout: (100 * 2.0 + 100 * 2.0) = $ 400, and the bet odds in the "Parlay" are equal to 400/200 = 2.0.Bavaria win with a difference of 1 goal: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 0.0 + 100 * 1.0) = $ 100, and the bet odds in the "Parlay" are equal to 100/200 = 0.5.Bavaria lose or draw: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.If the "double Asian handicap" bet of $ 200 was placed on: Juventus
Juventus wins or draws: both "halves" of an initial "double Asian handicap" bet win (+1 and +1.5).
Payout: (100 * 1.8 + 100 * 1.8) = $ 360, and the bet odds in the "Parlay" are equal to 360/200 = 1.8.Juventus loses with a difference of 1 goal: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 1.8 + 100 * 1.0) = $ 280, and the bet odds in the "Parlay" are equal to 280/200 = 1.4.Juventus loses with a difference of 2 goals or more: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.Example 4 Team Double Asian Handicap Odds Bavaria: -1.75 (-1.5, -2) 2.0 Juventus: +1.75 (+1.5, +2)) 1.8 If the "double Asian handicap" bet of $ 200 was placed on: Bavaria Madrid
Bavaria win with a difference of 3 goals or more: both "halves" of the initial "double Asian handicap" bet win (-1.5 and -2).
Payout: (100 * 2.0 + 100 * 2.0) = $ 400, and the bet odds in the "Parlay" are equal to 400/200 = 2.0.Bavaria win with a difference of 2 goals: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 2.0 + 100 * 1.0) = $ 300, and the bet odds in the "Parlay" are equal to 300/200 = 1.5.Bavaria lose or draw: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.If the "double Asian handicap" bet of $ 200 was placed on: Juventus
Juventus wins or draws: both "halves" of an initial "double Asian handicap" bet win (+1 and +1.5).
Payout: (100 * 1.8 + 100 * 1.8) = $ 360, and the bet odds in the "Parlay" are equal to 360/200 = 1.8.Juventus loses with a difference of 1 goal: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 0.0 + 100 * 1.0) = $ 100, and the bet odds in the "Parlay" are equal to 100/200 = 0.5Juventus loses with a difference of 2 goals or more: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.Example 5 Team Double Asian Handicap Odds Bavaria: -2.25 (-2, -2.5) 2.0 Juventus: +2.25 (+2, +2.5) 1.8 If the "double Asian handicap" bet of $ 200 was placed on: Bavaria Madrid
Bavaria win with a difference of 3 goals or more: both "halves" of the initial "double Asian handicap" bet win (-2 and -2.5).
Payout: (100 * 2.0 + 100 * 2.0) = $ 400, and the bet odds in the "Parlay" are equal to 400/200 = 2.0.Bavaria win with a difference of 2 goals: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 0.0 + 100 * 1.0) = $ 100, and the bet odds in the "Parlay" are equal to 100/200 = 0.5.Bavaria win with a difference of 1 goal: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.If the "double Asian handicap" bet of $ 200 was placed on: Juventus
Juventus win or draw or lose with a difference of 1 goal: both "halves" of an initial "double Asian handicap" bet win (+2 and +2.5)
Payout: (100 * 1.8 + 100 * 1.8) = $ 360, and the bet odds in the "Parlay" are equal to 360/200 = 1.8.Juventus loses with a difference of 2 goals: you lose one "half" of the initial "double Asian handicap" bet and the other "half" bet has the result of a "return".
Payout: (100 * 1.8 + 100 * 1.0) = $ 280, and the bet odds in the "Parlay" are equal to 280/200 = 1.4Juventus loses with a difference of 3 goals or more: both "halves" of the initial "double Asian handicap" bet lose.
Payout: (100 * 0.0 + 100 * 0.0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0.0.2Example of double total bets # Match Double Total Bet Odds Bet 1 Milan-Juventus Over 2.25 (2, 2.5) 2.0 $200 2 Under 2.25 (2, 2.5) 1.8 $200 3 Over 2.75 (2.5, 3) 1.9 $200 4 Under 2.75 (2.5, 3) 1.9 $200 For example, if 2 goals were scored in the match, then we get these results in bets settlement:
Bet 1. First "half" bet (Over 2) - refund, the second "half" bet (Over 2.5) - lose. Payment: (100*1.0+100*0) = $ 100, and the bet odds in the "Parlay" are equal to 100 / 200 = 0.5
Bet 2. First "half" bet (Under 2) - refund, the second "half" bet (Under 2.5) - lose. Payment: (100 * 1.0 + 100 * 1.8) = $ 280, and the bet odds in the "Parlay" are equal to 280/200 = 1.4
Bet 3. First "half" bet (Over 2.5) - lose, the second "half" bet (Over 3) - lose. Payment: (100 * 0 + 100 * 0) = $ 0, and the bet odds in the "Parlay" are equal to 0/200 = 0
Bet 4. First "half" bet (Under 2.5) - win, the second "half" bet (Under 3) - win. Payment: (100 * 1.9 + 100 * 1.9) = $ 380, and the bet odds in the "Parlay" are equal to 380/200 = 1.9
- We reserve the right to offer other forms of wagering.
- The bets are accepted based on the line, i.e. the list of upcoming events at the odds for their outcomes offered by us.
- Line changes (handicaps, winner odds, totals, limitations for parlays, maximum bet amounts, and others) can happen after any bet, but the terms for already placed bets remain without change. The bets are accepted for the amount that does not exceed the account balance.
- The amount of bets for any event during one betting day cannot exceed the maximum bet for the event shown on the line (for the line of the same date). If there are bets with a total amount exceeding the maximum bet, such bets will not be accepted.
- The bets are accepted before the event starts. The bets that are placed after the start of the event are not valid and such bets will be refunded. These bets are removed from parlays. The exceptions are so-called live bets during the event. These bets are valid and will be refunded only if for some reason the bets have been placed after the event has ended.
- The results update in the progress of live events is for information purposes only. The result listed incorrectly cannot be the ground for canceling the bet. We are not responsible for the correctness of results during live events.
- The date and the time for the event start specified in the line are informative only. A wrong date is no reason to refund the bet if the bet has been placed before the start of the event. For bets settlement, the actual time of the event start is used. This time is determined according to the official documents (websites referenced by us when collecting information) of the organization that holds the competition, match, etc.
Football
Bets on football are accepted on the regulation time, which is conditioned by the rules of the match/tournament. It includes additional or compensated time appointed by the referee or by a representative of the match at the expiration of the regulation time (match/half). The minutes of the events that took place at the time compensated/added by the judge are considered to be 45 (forty fifth) or 90 (ninetieth) minute.
- The football bets are available for one team to win or draw or with handicap.
- Various values for handicaps and totals are offered for betting.
- Wagering on the total number of goals to be scored (total). If the number of goals falls exactly on the spread, the bets will be refunded.
- In a cup competition that consists of many matches, a wager called Qualifying to the next round is offered. The outcome of this bet is determined by the combined score of the relevant matches.
- "Who will score more goals" betting. It is proposed to compare two teams from different matches. The team that scores more goals wins. In the case of equal performance, the bet is settled at the odds 1. "Who will score more goals" bets are not accepted in parlays and systems that include bets with the same team.
-
You can wager on the individual total of each team (at the odds on the list). If the amount of goals scored falls exactly on the spread, the bets will be refunded.
You can wager on the double result:- 1X - win for home team or draw;
- Х2 - win for away team or draw;
- 12 - win for home or away team.
-
You can bet on the progress of the match:
- wager "match result" - on the outcome of the match;
-
wager "halftime - fulltime" - on the outcome of the match at halftime and at full time.
To designate these outcomes in the line, the initial letters are used: W - win, D - draw. The outcome of the 1st half is indicated on the first place, and the match result - on the second. The outcome of the 2nd half is irrelevant for the bet settlement.
Example:
If the match ended with the score 1:1, and the first half - 1:0, then "W1D" outcome wins
If the match ended with the score 1:0, and the first half - 0:0, then "DW1" outcome wins
If the match ended with the score 1:0, and the first half - 1:0, then "W1W1" outcome wins - "goal in the first half" - wagering if there will be a goal scored in the 1st half "yes-no";
- wagers on the time of the 1st goal, the team to score first, total of the 1st and 2nd period and other bets that can be seen on the daily line.
- bets on statistics for a player in the match (the number of goals, fouls, yellow and red cards, etc.) are only valid if the player goes in starting lineup. A goal counts only if a player scored in his opponent's net. Own goals are not counted.
- "At least one player will score 2 or more goals" betting. Goals scored by substitutes also count.
- Wager home-away. Wagering on the home or away teams to cover the handicap.
-
Wager "round total". It is proposed to predict whether the total amount of goals scored by the home or away team will be over or under the proposed number. You can place wagers on the following events on the playing round:
- number of "home" wins;
- number of draws in the round;
- number of "away" wins;
- number of away team points, etc.
- If one of more matches are postponed and not finished within 12 hours, the bets for "home-away" and "round total" wagers will be settled at the odds 1.
- Betting forms "home-away" and "round total" are not allowed in parlays with matches of the same round.
-
The betting forms "number of corners, yellow cards, shots on target, offsides, substitutes, ball possession % and fouls will be offered with handicaps and totals (and individual totals).
Second yellow card that results in the player's sending off is counted as both: another yellow card and a red card. A player will be considered as sent-off if he was in the field at the time of being sent off. The cards shown to players on the bench and the staff (coaches, managers, etc.) are not included. Cards shown after the whistle about the end of the 1st half are considered as cards shown in the 2nd half. If two or more players from different teams get yellow/red cards in the match on the basis of a single incident of the game, then bets on the market "First yellow/red card" will be settled at the odds 1.
Corner kicks that are assigned but not performed do not count. If the referee does not count the corner and appoints it again - it will be counted only once. - Wagering "When the first (last) goal will be scored" with providing time intervals in the match. It is proposed to specify in what time interval the first (last) goal in the match will be scored. In the case of a 0-0 draw - settlement at the odds 1.
- "Who will score the next goal" betting. There are three possible outcomes: "Team 1", "Team 2", "No one will score."
- Games: "The most productive match", "The least productive match", "The match with the highest (lowest) number of goals". You are asked to select the one of the offered matches in which the highest/lowest number of goals will be scored. If the highest number of goals is scored in two or more matches, bets on these matches are settled at the odds 1. If one or more of these matches have been postponed for more than 36 hours from the time specified in the line, or were interrupted and not completed within 12 hours, all bets on these games will be settled at the odds 1.
-
The correctness of results can be checked from the following sites:
-
England. Premier League www.premierleague.com/
-
Germany. Bundesliga www.bundesliga.de/de/liga/
-
France. Ligue 1 www.lequipe.fr/Football/Ligue-1/
-
Italy. Serie A www.raisport.rai.it/
-
Spain. Primera Division www.lfp.es/
-
Europa League. Champions League www.uefa.com
If there is no information on these sources or there are obvious errors in reports, we reserve the right to carry out the settlement on the basis of internal sources.
-
-
Betting on the event that happens first in a definite period of the game. It is proposed to guess what will happen earlier in this period: out, foul, goal kick, corner, offside, goal or none of the above.
We also offer the pairwise comparison of various statistical indicators, in which it is proposed to find out what happens earlier at a certain period of the game. For example: out, foul or none of the above. Events are considered to have happened for the settlement purposes:- corner - in the case of actually performed kick from the corner mark;
- goal kick - when actually performed kick from the goal;
- out - in the case of actually performed throwing;
- offside - in the case of actually performed free kick after registering offside;
- foul - in the case of actually performed penalty or a free kick after registering violation;
- Time tracking of each event for the settlement purposes shall be considered:
- out - the time at the moment of actual throwing performance;
- foul - the time at the moment of actually performed penalty or a free kick after registering violation;
- goal kick - the time when actually performed kick from the goal;
- corner - time at the moment of the actual kick performance from the corner mark;
- offside - the time at the moment of actual free kick performance after registering offside;
- goal - time of the timer in the moment of scoring (when the ball crossed the goal line).
The settlement will be made on the basis of the timer on the TV broadcast of the TV channel, which will be specified in the comments to a specific match on the line. The line timer is started in synchronization with the specified channel timer and is considered the main timer for settlement in the moments of the temporary absence of a timer in a television broadcast. In case of any disputes, to make the final decision, information about the match we receive from internal sources will be applied.
-
What will happen earlier in the first 5 minutes of the match?
The rule applies to the bets placed before kick-off.
It is proposed to specify the event that will happen earlier within the first 5 minutes of the match:- Goal scored
- Corner appointed
- Penalty appointed
- Card shown
- None of the above
To consider the bet as the winning, the goal must be scored within the allotted time period. The goal will be ignored if the kick was made within the allotted time period, but the goal was scored out of this period.
Corner should be appointed within the allotted time period. Corner will be accepted if it has been appointed.
Penalty should be appointed within the allotted time period. Penalties will be ignored if the penalty happened within the allotted time period, and its appointment - no. Penalty will be considered if it has been appointed within the allotted time period, but its direct attempt to implement - not.
Red and yellow cards are counted at the moment when they are shown, and not at the time of the violation. The card is not considered if its cause happened within the allotted time period, but its immediate showing - no. Only those cards count that have been shown to the players participating in the game at the moment of penalty.
Appointed penalty is considered as the winning outcome if the penalty had been appointed before the referee showed the card for a foul. -
European handicap.
It is proposed to specify the outcome of the match based on the European handicap. Within betting on European handicap, game begins conditionally from a certain score, specified in the value of the handicap. The final result for these bets is settled by adding up the values of the score and the handicap.
European handicap does not mean the bet partial or full refund. The player can either win or lose the whole amount of the bet.
Possible outcomes in the European Handicap betting:- Home win with handicap.
- Draw with handicap
- Away win with handicap.
Example 1. Handicap (0-2) - a bet that suggests that after the end of the match we must add "0" to the Home scored goals, and "2" - for Away scored goals. Then the bet will be settled.
The outcome "Home win with handicap" will be winning if the match ends in a home win in 3 or more goals. For example: 3-0 (3-2 considering the selected handicap 0-2), 4-1 (4-3 considering the selected handicap 0-2), etc.
The outcome "Draw with handicap" will be winning only for Home team win in exactly 2 goals. For example: 2-0 (2-2 considering the selected handicap 0-2), 3-1 (3-3 considering the selected handicap 0-2), etc.
The outcome "Away win with handicap" will be winning if the Away team will not lose with a difference of more than 1 goal, will win or play a draw. For example: 1-0 (1-2 considering the selected handicap 0-2), 2-2 (2-4 considering the selected handicap 0-2), 1-2 (1-4 considering the selected handicap 0-2), etc.
Example 2. Handicap (1-0) – a bet that assumes that after the end of the match we must add "1" to the Home scored goals, and "0" - for Away scored goals.
Then the bet will be settled.
The outcome "Home win with handicap" will be winning for any win or a draw. For example: 1-0 (2-0 considering the selected handicap 1-0), 2-2 (3-2 considering the selected handicap 1-0), etc.
The outcome "Draw with handicap" will be winning only for Away team win in exactly 1 goal. For example: 0-1 (1-1 considering the selected handicap 1-0), 2-3 (3-3 considering the selected handicap 1-0), etc.
The outcome "Away win with handicap" will be winning only for Away team win in 2 or more goals. For example: 0-2 (1-2 considering the selected handicap 1-0), 1-4 (2-4 considering the selected handicap 1-0), etc. -
Shots on the goal post or crossbar.
The bet settlement takes into account shots at the goal post or crossbar, after which the ball stayed in play (touched the player, referee on the field, another goal post or crossbar, went on the pitch or crossed sideline. A goal post or crossbar is not taken into account if a) the game was stopped before the ball touched the goal post or crossbar b) after the ball touched a goal post or crossbar, it went out of the goal-line without touching the field, c) ball flew into the net (the goal was scored). -
Goalless time intervals.
If there are n goals scored in the match, the minutes on which the goals were scored break the time interval of the match (from 0 to 90 minutes) to n + 1 goalless time intervals. The length of the goalless time intervals is settled as the difference between the minutes of the first goal and the beginning of the match, the minutes of successively scored goals and the minutes of the end of the match and the last goal. The beginning of the match - 0 min. If there is no goals in match, we believe that in the match one goalless time intervals with a length of 90 minutes.
Example 1. If the match in regular time was finished with a score of 0:0, so there was only 1 goalless time interval with 90-minute length in the match.
Example 2. If there are 2 scored goals in match: 23 and 62 minutes. In this case, there were only 3 goalless time intervals in the match: 23 minutes (23-0), 39 minutes (62-23) and 28 minutes (90-62).
Example 3. If there are 3 scored goals in match: 15, 45 and 90 minutes. In this case, there were only 4 goalless time intervals in the match: 15 minutes (15-0), 30 minutes (45-15), 45 minutes (90-45) and interval with 0 minutes (90-90). -
Cyber Football
Bets are accepted on a computer game FIFA (football simulator with 2 players take part in).
Matches are played with game settings “2 halves by 5 minutes each ” complexity level “beginner”.
If the match is stopped due to technical reasons (computer failure, disconnection, etc.) and re-match appointment, bets on stopped game will be settled according to main rules of settling interrupted events.
Bets on re-match is accepted as for a new event.
If due to technical failure the live streaming of the event is interrupted (disconnection, DDoS, etc.), this is not a reason for canceling bets, except for situations in which it is impossible to know the final result of a match.
Ice hockey
- The ice hockey bets are available for one team to win or draw or with handicap.
- Wagering on the amount of goals to be scored (total). If the amount of goals falls exactly on the spread, the bets will be refunded.
- We offer a variety of handicaps and totals.
- Wagers/bets are accepted only for the regular time of the matches (no extra-time). However, preliminary predefined cases may apply.
- Wager on the productivity of each period. It is proposed to guess which period will have more goals scored.
- You can wager on the individual total of each team. If the amount of goals scored falls exactly on the spread, the bets will be refunded.
-
You can bet on the progress of the match:
- win of one team or draw or win with handicap, offered for a predetermined period;
- number of goals scored in a predetermined period;
- wagering on the time of first goal, first team to score and other forms of betting;
- wagering on how many points a predetermined player will score, goals and assists will be added together in a specific match;
- Wagering on shots on goal, penalty time and individual performance of players are accepted including overtime period, excluding shootout (for NHL and World Cup matches). For the matches held under the auspices of the International Ice Hockey Federation (IIHF), bets on shots on goal, penalty time and individual performance of players are accepted for the main time
- Faceoffs bets are accepted only for Faceoffs won (marked as FW), except special cases.
-
Wagering on the betting form "home-away".
- wagering on the difference of goals scored by home and away teams compared to the handicap;
- wagering on the amount of total goals scored by home and away teams compared to handicap;
- wagers on the following events on the playing round (how many goals will home teams score, how many goals will away teams score, how many home wins, how many draws, how many away wins, how many points will home teams get, how many points will away teams get etc.)
- If one or more matches are interrupted and not finished, the bets on the "home-away" will be settled at the odds 1.
- The wager "home-away" is not accepted in parlays with other wagering events of the same round.
- Wager "Come From Behind and Win". In this bet type the team you have betted on must be losing in some point in the match and then must come back and win the game.
- Wagering "When will be scored the first (last) goal" with providing time intervals in the match. It is proposed to specify in what time interval the first (last) goal in the match will be scored. In the case of a 0-0 draw - settlement at the odds 1.
-
"Who will score the next goal" betting. There are three possible outcomes:
"Team 1", "Team 2", "No one will score." - If an ice hockey match is not played for some reason on the same day as mentioned on the list, the bets on that match will be refunded.
-
Cyberhockey
- Bets are accepted on a computer game NHL (hockey simulator with 2 players take part in).
- Matches are played with game settings “3 periods by 3 minutes each ”, difficulty level of the game and goalkeeper difficulty level - “Semi-Pro”, game style – “Arcade”.
- If the match is stopped due to technical reasons (computer failure, disconnection, etc.) and re-match appointment, bets on stopped game will be settled according to main rules of settling interrupted events.
- Bets on re-match is accepted as for a new event.
- If due to technical failure the live streaming of the event is interrupted (disconnection, DDoS, etc.), this is not a reason for canceling bets, except for situations in which it is impossible to know the final result of a match.
-
You can check the correctness of the results on the following websites:
-
www.nhl.com NHL
-
www.khl.ru KHL
-
www.vhlru.ru VHL
-
www.mhl.khl.ru MHL
-
www.iihf.com World Championship
-
www.championshockeyleague.net Champions Hockey League
-
www.hockey.by Belarus
-
www.ishockey.dk Denmark
-
www.deb-online.de Germany
-
www.hockey.no Norway
-
www.hokej.sk Slovakia
-
www.finhockey.fi Finland
-
www.hokej.cz Czech Republic
-
www.sehv.ch Switzerland
-
www.swehockey.se Sweden
-
Tennis
-
The following wagers are offered on tennis:
- winner of the match;
- correct score in sets;
- handicap for games in the match, set;
- handicap for amount of sets in match;
- amount of games and sets played in match;
- amount of aces scored by a player in match;
- amount of double faults for a player in match;
- winner of the set;
- whether there will be a tiebreak or not (score 7:6 or 6:7).
Possible Live bets:- Who will win the n point in the game (you have to differentiate between the number of points in the game and score of the game. For example, if the score of the game is 15:15, then 2 points have played in a game, but not 30. Also, the 3rd point is being played out now, and the 4th point will be played next);
- whether there will be the score 40:40 in a game (it means that during the game the score will be 40:40);
- score of the game (in this case W1:15 means that the 1st tennis player will win the game, and the second one will win only 1 point in a game; 40:W2 means that the 2nd tennis player will win the game after the score of 40:40), etc.;
- Who will win the first n games in the set, etc.;
- The company uses its own data about the actual course of the game for the live bets settlement. If for any reason (such as loss of broadcasting, the lack of results in the information sources, etc.) the outcome of events cannot be determined, these events will be settled at the odds 1. Claims for bets placed on a live event, are accepted not later than 24 hours after the end of the event.
- All bets on the match are valid until the end of the tournament.
- Information on the tennis court surface is informative only. If the surface is changed, wagers will remain valid. Claims regarding the incorrect information about the court surface in our line are not accepted.
-
If one of the players in tennis refuses to continue the match regardless of the reason (or disqualified), then:
-
if at least one set is completely played:
- all bets on winning the match are considered valid, and the match winner is determined in accordance with the decision of the referee;
- all bets that are actually completed by the time of interruption are considered valid;
- all bets on handicap for the games and sets will be refunded regardless of the score.
-
if the first set is not completely played:
- all bets on the match will be refunded, with the exception of actually completed by the time of match interruption.
The match is interrupted at the score (6:4 4:4).
Bet on the 1st set, "Djokovic handicap +1.5" - lost;
Bet "Total of 2nd set over 7.5" - won;
Bet "Total of 2nd set under 8.5" - return;
Bet "Total of 2nd set over 9.5" - return;
Bet "Total of match over 19.5" - return;
Bet "Total of match under 17.5" - loss (because actually 18 games have been already played).
Example 2. The match Nadal - Djokovic.
The match is interrupted in the first set with the score 5:5.
Bet on the 1st set, "Djokovic handicap + 2.5" - return;
Bet on the 1st Set "Nadal handicap - 2.5" - return;
Bet "Total of the 1st set over 9.5" - won;
Bet "Total of the 1st set over 10.5" - return;
Bet "Total of the 1st set under 10.5" - return;
Bet "first to win 4 games" - settlement in accordance with the result. -
if at least one set is completely played:
-
If the regulations are changed (a 5-set match is changed into a 3-set match or vice versa), the bets on the following:
- winner of the match;
- winner of the 1st set;
- handicap for games in the 1st set;
- total of the 1st set;
- will be settled at the odds offered in the line, other wagers will be refunded.
- If in the team competition one or more players are changed in the team, no matter what the reason, the bets on the match outcome and qualifying to the next round will stand. If the players have been nominated on list in doubles and one of the players is changed, the odds will be settled as "1". If the players have not been nominated on the list, the bets will stand.
- If the 1st place is shared between two or more winners, the bets on these players to win the tournament will be settled at the odds 1.
- If instead of the decisive set of the match, the so-called "champions tie-break" is played (short game to 10 points won), then the score of the 3rd set is settled as 1-0 for games. Accordingly, the overall Total and handicap for games in a match are settled on the basis of 1-0 in the 3rd set.
American football
- In American football, you can wager on one of the teams to win or win with handicap. In case of a tie, the bets will be refunded.
- Wagering on total (how many points will be scored) is offered. If the total falls on the spread, the bets will be refunded.
- Buying is not allowed in American football.
- If for some reason the match was not played on the announced day, the bets will be refunded.
Baseball
- Bets on totals and handicaps are action if at least 7 innings are completed (or 6,5 innings if the team batting second is ahead). If the game is postponed or cancelled while the extra-innings, the final score is determined according to the score after the last full inning (unless the team batting second scores to tie, or takes the lead in the bottom half of the inning). Thereby, the winner is determined according to the current score at the moment of the game interruption.
- Bets for the 5 innings outcome are calculated, based on the result after 5 full innings.
- Minimal duration of a game for bets to have action is 5 full innings (unless the team betting second is ahead after 4,5 innings).
- Regardless of the starting pitcher change, all bets are action and replacement has no influence on the calculation.
- If the game did not take place at the time it was originally scheduled and has been postponed for more than 36 hours from the time indicated in the coupon, all bets are void and calculated with the odds “1”.
- If a game was interrupted and postponed for more than 12 hours from the time indicated in the coupon, the winner has to be determined according to the score at the moment of the interruption (unless the team batting second scores to tie, or takes the lead in the bottom half of the inning). Thereby, the winner is determined according to the current score at the moment of game interruption. If a game interrupts after home team tied, all bets are void and calculated with the odds “1".
- If a game is not completed because of any reasons (rain, weather conditions or other external factors), the winner has to be determined according to the score at the moment of the interruption, and this score confirmed as a final result by the official site of the tournament. If the official site of the tournament does not confirm this score as a final result, all bets are void and calculated with the odds “1”.
- If a game ended in a draw, all bets are void and calculated with the odds “1".
Basketball
-
Bets on basketball fixtures are accepted with overtime included in case of a draw in the scheduled game time except for the bets on Wins and a Draw (1, X, 2) that are accepted in the scheduled game time only. In basketball fixtures, where the rules of the tournament for determining the winner of the confrontation provided for two games (at home and away), the match can end in a draw without overtime. Overtime, appointed by the sum of the results of the two matches, is taken into account only in the betting rounds for the next round (Qualifying to the next round).
Example 1: match play-offs (according to the rules of the tournament, the winner of the confrontation is determined by the sum of two meetings), the first match ends with a score of 79-71 (without OT), the second match - 80-80 (without OT). In the second match, overtime is NOT stipulated by the regulation, since the score for the sum of two meetings is 159 - 151, as a result of qualifying to the next round of the first team.
Example 2: match play-off (according to the tournament rules, the winner of the confrontation is determined by the sum of two meetings), the first match ends with a score of 85-80 (without OT), the second match - 96-91 (without OT). On the sum of two meetings 176 - 176. In the second match, Overtime is appointed to determine the team that will qualify to the next round (this OT is not taken into account when settling the bets on the match, except for the outcome Qualifying to the next round). - Wagering is offered on totals (combined and individual), either on the combined score of both teams or the score of one individual team.
- In some specific matches, wagering on quarters is offered. In addition, the bets can be placed on the most/least productive quarter.
- The number of points, rebounds scored by the players of either team according to the line. In case the player does not enter the field, the bet will be refunded.
- "Duel of players". In this wagering, it is proposed to guess which player will score more points, assists, or rebounds. If one of the players did not participate in the match, all the bets will be refunded. Bets are settled according to official records (including overtime period).
- The "Performance" market counts only points, rebounds and assists of a certain player. Blocked shots and steals are not counted as Performance.Bets are settled according to official records (including overtime period).
Volleyball
- Bets are settled according to the accepted odds and all other bets will be refunded.
- Handicaps and totals for the match are specified in points, unless otherwise indicated in the line.
- 1-4 sets in volleyball match are played to 25 points, unless otherwise indicated.
-
The following wagers are offered on volleyball:
- to win the match or set;
- Total of a match or set;
- handicap in the match or in the set;
- whether there will be the tiebreaker in the match (tiebreaker in the volleyball match is called the fifth set).
-
If regulations of the match are changed (1-4 sets are played to 21 points, or vice versa), then bets on:
- winning team in a match;
- the victory of the team in the set;
- exact score of the match;
- “Golden set” is not taken into account to settle the match.
Beach Volleyball, Table Tennis, Badminton
- Handicaps and totals for the match are specified in points, unless otherwise noted.
-
The following wagers are offered on volleyball:
- to win the match or set;
- total of a match or set;
- handicap in the match or in the set;
- whether there will be the tiebreaker in the match (a tiebreaker in the volleyball match is called the fifth set).
-
If the match is interrupted or one of the teams for any reason refuses to continue the game or is disqualified, then:
-
if at least one set is fully completed:
- all bets to win the match are considered valid, the winner of the match is determined in accordance with the referee's decision;
- all bets actually completed by the time of the match interruption are considered valid;
- all bets on the games and sets handicaps are subject to refund regardless of the score.
-
if the first set is not fully completed:
-
all bets on outcomes, the results of which are determined before the time of match interruption are settled by these results. Bets on outcomes which results are not determined are settled with odds of 1.
Example of beach volleyball bets settlement in situations when the second team refuses to continue due to injury: The match was interrupted in the second set when the score was 1:0 (21:15, 19:20): - bet "2nd set total over (38.5)" - win
- bet "2nd set total under (38.5)" - loss
- bet "2nd set total over (40.5)" - refund
- bet "second team to win" - loss
- bet "first team's handicap in the match (-5.5)" - refund
-
all bets on outcomes, the results of which are determined before the time of match interruption are settled by these results. Bets on outcomes which results are not determined are settled with odds of 1.
-
if at least one set is fully completed:
Water Polo, Handball, Beach Soccer, Floorball, Futsal, Field Hockey
Wagering on these kinds of sport is for regular time only due to the rules of the match, unless otherwise specified.
Formula 1, Nascar, MotoGP
-
The following wagers are offered:
- Wagering on the winner of the race. The driver who takes the first place in final classification is the winner.
- Bet on a certain place of the driver in the final protocol.
- Wager "Finish - Not finish". Prediction whether a driver will finish the race or not. If a driver is qualified, he is considered to finish the race.
- Wager "Fastest lap". Which driver will drive the fastest lap.
- Wager "Who is higher". You predict which driver in a predetermined pair will finish in a higher position after the race. If both drivers fall out of the race, the driver that finished more laps is the winner. If both drivers fall out of the race on the same lap, the bets will be refunded.
- Wager "Which team is better in a race". The predetermined teams are put in pairs consisting of two drivers and you predict which team will perform better. The better team will be determined by: a) which team has more drivers to finish the race and be classified; b) if both teams have the same amount of classified drivers, the better team is determined by adding up the sum of drivers` positions. The smaller sum, the better a team has performed. (If the sum is the same for both teams, the bets will be refunded). If drivers are changed in teams after the line has come out, all bets on these teams will be refunded.
- The positions of the drivers are determined by the official final protocol published immediately after the race. Subsequent disqualifications of the drivers and the changes in the protocol after the race will not be considered.
- Warm-up lap is included in the race result.
- In the disputable situations regarding winner bets and podium bets, the results will be determined according to the price giving ceremony.
- All bets are valid until the end of the stage, regardless of the possible transfer of the start time and date.
-
We also offer the following types of qualification bets:
- Bet on the winner of the qualification. The winner is the driver who has won first place in the specified qualifications.
- Bet "Who is higher". You predict which driver in a predetermined pair will finish in a higher position after the qualification. The driver who showed the best qualification time in qualifying lap and took a higher place on the protocol of the specified qualifications is considered higher.
- Bets on the driver who did not show qualification time in the specified qualifying lap as well as bets on comparison of him to other will be refunded.
- The correctness of results can be checked from the following sites:
UFC
- The following types of bets are offered:
- In the case of a draw at UFC tournaments, bets on winning of the participants are settled at the odds "1". The remaining outcomes are settled according to the actual result.
- Win - includes winning by Decision (DEC), winning by KO, TKO, winning by Technical Decision (TD) based on points, Disqualification or Failure, winning by submission or surrender of an opponent.
- Ahead-of-time win – includes winning by KO, TKO, Disqualification or Failure, winning by submission or surrender of an opponent
- The duration of the fight. Predict a specified number of rounds. Half of the 5 minute round considered 2 minutes. 30 sec. Therefore, 2.5 rounds - this will be 2 minutes 30 seconds of the 3rd round. If the match ends exactly in the middle (2 minutes 30 seconds) of the round, the over/under rate on the total number of rounds held is returned to the players.
- Win in the round - if the fighter refuses to continue the fight in the break between rounds, or after hitting the gong about the beginning of the next round, does not go out, then the last round held is considered to be winning.
- If for any reason the selected number of rounds changes, then all bets on number of rounds will be refunded. The bets on the result of the fight and the method of its achieving remain in force.
CyberUFC
- Bets are accepted on a computer game UFC (fighting simulator with 2 players take part in).
- Matches are played with game settings “3 rounds by 3 minutes each”, complexity level of the game - «Legend», match mode – “accelerated”, fighting in the ground (wrestling) and painful techniques without prompting. vIf the match is stopped due to technical reasons (computer failure, disconnection, etc.) and re-match appointment, bets on stopped game will be settled according to main rules of settling interrupted events.
- Bets on re-match is accepted as for a new event.
- If due to technical failure the live streaming of the event is interrupted (disconnection, DDoS, etc.), this is not a reason for canceling bets, except for situations in which it is impossible to know the final result of a match.
Boxing/Mixed Martial Arts
The following types of bets are offered:
- Win - includes winning by Decision (DEC), winning by KO, TKO, winning by Technical Decision (TD) based on points, Disqualification or Failure, winning by submission or surrender of an opponent.
- Ahead-of-time win – includes winning by KO, TKO, Disqualification or Failure, winning by submission or surrender of an opponent.
- The duration of the fight. Predict a specified number of rounds. Only the number of complete rounds counts for settlement purposes.
- Victory in the round - if the fighter refuses to continue the fight in the break between rounds, or after hitting the gong, about the beginning of the next round, does not go to him, then the last round that is held is considered to be winning.
- If for any reason the selected number of rounds changes, then all bets on number of rounds will be refunded. The bets on the result of the fight and the method of its achieving remain in force.
- In the case of a draw in mixed martial arts, bets on winning of the participants are settled at the odds "1". The remaining outcomes are settled according to the actual result.
Australian Football
- You can bet on the victory of one of the teams that compete in a match (including bets on the victory with handicap). Bets on a match are accepted without overtime, unless otherwise specified in the line. In case of a draw, all the bets on the victory will be refunded.
- You can also bet on "Total" (the total number of points scored in the match). If the bet is equal to "Total", it will be refunded.
- Buying points is possible for bets on "Total" or "Handicap".
- We also offer bets on separate quarters and halves of matches.
Bandy
- The wagers on bandy are accepted only for regular time (except for the cases set out specifically).
- Various values of handicap and total are offered for betting.
- Wagering on individual totals of teams is offered (if odds are available). If the score falls on the spread, the bets will be refunded.
- Wagering on the points scored in each period is not available.
- If for some reason the match is not finished in the regular time, the bets will be refunded.
Biathlon
- Wagering on winner. Winner is determined by the official protocol. If there is more than one winner - bets on these participants will be settled at the odds 1. If one of the participants mentioned in the list does not start - bets on him will be settled at the odds 1.
- Wager "Who is higher". You predict which competitor in a predetermined pair will finish at a higher position after the competition (team in relays). If both competitors fall out of the race or one of them does not start - the bets will be refunded.
- Bet on the determining the participant's finishing position. It is proposed to specify the place at which the participant will finish the race. Variants: 1-3 places, 4-10 places, 11+ (i.e. below the 11th place in final protocol).
- Wagering on the total number of target misses of the race participants. If the number of target misses is equal to handicap, all bets will be refunded. The bets that have actually finished by the time the participant falls out of the race (or disqualified) are considered valid.
-
Wagering on the number of misses in each shooting round (1st round, 2nd, 3rd, 4th, etc.). The bets that have actually finished by the time the participant falls out of the race (or disqualified) are considered valid.
The participant has fallen out of the race after two shooting rounds. Bets on the number of misses in the 1st and 2nd shooting rounds are valid. Bets on the 3rd and 4th shooting rounds will be settled at the odds 1. - Comparison - which participant will make more target misses in a predetermined pair. You should predict which of the two competitors will make more target misses. If one of the competitors does not start or finish the competition - bets will be refunded.
- Comparison "Member of which of the teams will be higher". Select a team whose member finishes higher on the final protocol.
- The final results of the race participants will be determined by the official protocol immediately after the competition. Possible changes after that like disqualification will be ignored.
- You can check the statistics on the official website
Pool, Darts, Snooker
-
The following types of wagers are offered:
- Winner of the match;
- Handicaps by games in the match;
- Total games in the match.
- All the bets on matches are valid until the end of the tournament in which they are held.
- If for some reason one of the participants refuses to continue the competition or he is disqualified, he loses all incomplete games, and the bet will be settled by the received result. If the refusal (disqualification) occurred before the beginning of the match, the bets will be refunded.
- If the match regulations have been changed, the bets placed on a winner of the match will be settled at the accepted odds while the bets placed on handicaps or totals will be refunded.
Rugby
- We offer to place bets on: outright winner of the match, draw, handicaps.
- Total number of scored points ("total"). If the number of scored points is equal to the total, the bets will be refunded.
- The bets are accepted only for regular time in case of wagering on winning or draw of either team. Possible overtime counts for handicap and "total" betting.
- If for some reason the match was not played on the announced day, the bets will be refunded.
Chess, Draughts/Checkers
- Wagering is offered on the winner of the match, draw, handicap, both for either participant and for the team.
- Qualification to the next round, winner of the tournament, "black-white".
Athletics
-
Two types of wagering are offered in athletics:
- Winner of the tournament: the winner is determined in accordance with the official final protocol that is issued immediately after competition ending.
- Who is better: the participant that has taken higher position in the final protocol is better. If both participants withdrew at the preliminary stage, a participant that has passed to a higher stage is considered to be better. If they have reached the same stage, the bet will be refunded.
- Final results of the participants of the competition are determined in accordance with the official final protocol that is issued immediately after competition ending. Further disqualifications and any changes in the final protocol are not taken into account by us.
- In the event of a player withdrawing before the start of competition, the bets will be refunded. In this case, the bet "winner – any other" will also be refunded.
- A participant is deemed to have started once he has reached any of the preliminary stages.
Golf
-
The following types of wagering are offered:
- Comparison of two participants of the tournament. The winner is the participant with the lowest score. In case of similar results, the bet will be refunded.
- Total of the 1st place score – the score of the player who was announced as a winner of the tournament.
- Total the 2nd place score – the score of the player who has finished on the second place in the tournament. If there are more than one participants who finished on the second place, score of only one participant's counts.
- Score of a player is the number of the ball strokes that were made while passing the holes during the tournament.
- In the event of a player withdrawing, the other opponent player is considered the winner. If both of the players withdraw, the lowest score after withdrawing will determine the winner.
- In the event of a player being disqualified after having teed off or withdrawing before the end of 2 rounds or before the end of the tournament, the opponent player is the winner.
- In the event of a player being disqualified during the third or fourth round and the opponent player has already withdrawn, the disqualified player is the winner.
- All the results are taken on the tournament official sites.
Cycling, Skiing (Race, Ski Jumping)
-
Such bets are offered:
- Race winner
- "Who is better": the participant that has taken higher position in the final protocol is better.
- Final results of the participants of the race are determined in accordance with the official protocol of the race that is issued immediately after its ending on the official site. Further disqualifications and any changes in the final protocol do not count.
- In case of a participant’s withdrawal from the race before its beginning, the bets will be refunded.
Curling
-
The following types of bets are offered:
- Winner of the match;
- Handicaps and total.
- Extra ends count for betting on match.
Bowls, Netball
-
The following types of bets are offered:
- Winner of the match;
- Handicaps and total (sets/quarters)
- If a victory in a match is awarded to a player/team before all sets/quarters are finished, then all bets on the match will be refunded, with the exception of actually completed by the moment of interruption.
Eurovision Contest
-
The following wagers are offered:
- Qualification to the final round and the winner;
- "Who is higher". You should compare the competitors in proposed pairs to select one of them to gain higher position in the final protocol. In case the results are equal, the bets will be refunded.
- Wagering on the position that a participant will gain in the final protocol.
- For details about the results of the contest, please see www.eurovision.tv
E-Sports
Features of the bets settlement:
- e-Sports bets are settled based on the General Rules, in view of specificity, which is inherent in a particular e-Sports discipline.
- e-Sports games in the format of Bo1, Bo2, Bo3, etc. (Best of 1, 2, 3, 5, etc.) - the total number of game maps, in which the selected team should reach the prevailing number of victories. The winner of the match is determined by the amount of won maps. For example: for Bo5 3 victories are required, for Bo3 - at least 2 maps, etc.
Dota 2 / League of Legends
- Final settlement is carried out according to data recorded immediately after the destruction of the main building (Fortress/Nexus) by one of the opponents. Similarly, the settlement will be carried out in the event that one of the teams has surrendered (castle/nexus in this case is not destroyed by the direct impact of the enemy). Victory is awarded to the opposing team.
-
2. In the case of the technical defeat in the round (on the map) or full match, outcomes that are not determined by gaming competition at the moment of the TD announcement are settled at the odds "1". Reasons of technical defeats: DDoS attack, lateness of a team or a participant, replacement of a player during the game and other circumstances by the decision of referees or organizers. Settlement at the odds "1" is possible only before the game start (exit of creeps / minions from the barracks on the line). When a technical defeat is appointed, or at any unplanned situations (disconnect, exit (leave) of players from either team, etc.), all further bets are settled according to the results.
A technical defeat on the map counts for bet settlement (win / Handicap / Total) for the full match. Numbering of maps in line in the case of the technical defeat also takes into account the round (map) that were won by TD. - This discipline provides totals and handicaps on the murder (kills) of the main characters and maps.
- The advantage of one map before the start of event. Under the decision of the referees / regulations, in some cases, the winning of one of the teams can be "in absentia" awarded in 1 map (do not confuse with a technical defeat or bet on handicaps). In outcomes with the total of rounds (maps) or winning in a particular (in sequence) map, this "absentia" map is ignored. That is, the first map in the line is considered to be the first map played by the participants.
- The "First Blood" (first kill) bet is interpreted as the murder of the game hero (champion) by the characters (champions) of the opposing team. The first murder in the map by neutral units or creeps / minions does not count and the bet stands until the first killing of a hero by characters of the opposing team. The "Will take first 10 kills on map" bet is offered for the team that will be the first to achieve 10 kills of the opposing team heroes.
- Re-host, rebuild, rollback of game situation to the moment of the autosave at a certain time (in the case of server crashes, network problems, DDOS-attacks, etc.) is not a reason for bets refunding.
Counter Strike: Global Offensive
- Winning the round is achieved by the murder of all rivals on the map, detonation / defusing the bomb or the end of the time on the timer of the round.
- The game is started after the first murder in the "pistol" round.
- Winning on the map is achieved by victory at least in 16 rounds (or, in accordance with the tournament regulations). In the case of a draw on the map (when the score by rounds is 15-15), the organizers may appoint additional 6 or 10 rounds, the so-called "overtime" (OT). Winning in overtime is awarded to the team that reaches the advantage with a gap of 2 rounds or, if one of the teams cannot "catch up" the opponent in the allotted number of rounds (for example: when the score is 19-15, 21-17 etc.). In case of a draw (21-21), referees appoint another 6 (10) additional rounds.
- If the event begins with the advantage in the score of one of the teams, according to the regulations, or by decision of the referee, all bets are settled at the odds "1", except for the cases when the information was indicated in line.
- In case of the match format change (the number of maps, rounds or other regulated standards), the bets are settled at the odds "1", except for the bets on the outcomes that have already been determined.
- Handicap and Total bets are offered both on rounds, maps and murder (kills) in a particular period of time.
- Bets on the outcomes that are actually determined by the time of the interruption / retire of teams or the technical defeat are settled based on the results. Bets, the result of which is not determined at the time of a stop, will be settled at the odds "1".
- Bets are accepted both on standard win / draw and the handicaps (kills, maps, rounds) as well as specific outcomes, such as: "The winner of the N round on the N map" or "The winner of the Tournament", etc.
Starcraft 2
- Bets on this discipline are presented in the form of winning outcomes as well as handicaps and totals.
- All bets are valid after the game starts, and the confirmation by the players and / or referees is received.
- In the event of the technical defeat assigned before the game, the bets are settled at the odds "1". In case of appointment of the technical defeat after the start of the game, all bets remain valid (provided that the outcome can be determined at the moment of interruption and the announcement of the technical defeat of one of the participants).
HearthStone
- Bets on this discipline are presented in the form of winning outcomes as well as handicaps and totals.
- All bets are valid after the game starts. The beginning of the game is when both players got the first card from the deck.
- In the case of a stop / interrupting the game (without the possibility to continue) or technical defeat, all outcomes that are possible to determine at this point are settled based on the results. The outcomes that are impossible to determine are settled at the odds "1" (assuming that the game will not be continued according to clause 37 hereof).
- The player who has not finished the game for any reason (DDoS attacks, network disconnection), loses in the current and subsequent games.
Overwatch
- Bets on this discipline are presented in the form of winning outcomes as well as handicaps and totals.
- All bets are valid after the game starts.
- In the case of a stop / interrupting the game (without the possibility to continue) or technical defeat, all outcomes that are possible to determine at this point are settled based on the results. The outcomes that are impossible to determine are settled at the odds "1" (assuming that the game will not be continued according to clause 37 hereof).
World of Tanks
- Bets on this discipline are presented in the form of winning outcomes as well as handicaps and totals.
- All bets are valid after the game starts.
- In the case of a stop / interrupting the game (without the possibility to continue) or technical defeat, all outcomes that are possible to determine at this point are settled based on the results. The outcomes that are impossible to determine are settled at the odds "1" (assuming that the game will not be continued according to clause 37 hereof).
PUBG
- Bets on this discipline are presented in the form of outcomes of head to head place match-up, total kills committed by both teams on a round, total kills committed by a player/team, handicap kills on a round, bets on getting into the top by a player or team that are represented in the line. Top - the number of surviving players/teams on the map.
- If the game for the player is interrupted due to his refusal - bets on all the markets of this game will be canceled, except when the results of the markets have already been determined at the time of stopping the game.
- If the game is interrupted for technical reasons on the server or technical problems of the player, bets on all markets of this game will be canceled, except when the results of the markets have already been determined at the time of stopping the game.
Dota 2. Auto chess
-
Bets on this discipline are presented in the form of outcomes for:
- Place in the top - position in the rating of the game which the streamer occupies by the time of the completion of his game.
- Total of the rounds played by the streamer until the end of his game. If 30 rounds have passed, and the player completes the game for 31, then the result of the total for the rounds is the number 31.
- Handicap - the difference between won and lost rounds.
- The game is considered as held at the start of the 5th round. If the game did not held (4 rounds were not played), bets are returned.
-
Any technical problems (loss of internet, crash, force majeure) entail a return of bets. Exceptions are cases like the following:
- 20 rounds were played, the bet was placed on a total of 15 over and a failure occurred. The total bet will be settled.
- 20 rounds were played, the bet was placed on a handicap +2.5 over and a failure occurred. Handicap bet +2.5 will be returned.
- The player entered the top 3, bet was placed at the entrance in the top 5 and crashed. The bet will be calculated if this entry into the top is guaranteed and getting out of the top 5 is impossible (example. the rest of the participants lost).
- The official data source is a stream record or the stream itself.
Cricket
-
To win/3-way betting:
- All match betting will be settled in accordance with official competition rules.
- In matches affected by adverse weather, bets will be settled according to the official result.
- If there is no official result, all bets will be void.
- In the case of a tie, if the official competition rules do not determine a winner then dead-heat rules will apply.
- In competitions where a bowl off or super over determines a winner, bets will be settled on the official result.
- In First Class Matches, if the official result is a tie, bets will be settled as a dead-heat between both teams and not as a draw. If the Draw was offered, bets on the Draw will be settled as losers in the case of a Tied match.
- If a match is abandoned due to external factors, then bets will be void unless a winner is declared based on the official competition rules.
- If a match is cancelled then all bets will be void if it is not replayed or restarted within 48 hours of its advertised start time. If there is a change in teams from originally advertised then all bets will be void.
- In First Class Matches, in case of Draw, To win(2-way) bets will be void.
-
Double Chance:
- A tie will be settled as a dead heat.
- All match betting will be settled in accordance with official competition rules.
- If there is no official result, all bets will be void.
-
Tied Match:
- All bets will be settled according to the official result at the end of the match not including super overs, bowl offs or any other means of deciding the winner in the event of a tied match. Must be an official match result for bets to stand.
- If the match is abandoned or there is no official result, all bets will be void.
- For First Class matches a tie is when the match has reached a natural conclusion with all innings being completed and the final scores are level. If all innings are not completed in their entirety at the end of the match then the official match result will stand. A declared innings counts as a completed innings.
-
Toss winner:
- If no toss takes place, all bets will be void.
-
Most Fours/ Most Sixes:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled in either innings due to external factors, including bad weather, unless settlement of the bet has already been determined. or the match has reaches its natural conclusion without either innings being shortened to less than 80% of the originally alloted overs. Bets will stand if either innings is shortened due to it reaching a natural conclusion.
- "In First Class matches resulting in a draw, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Only fours/sixes scored from the bat (off any delivery – legal or not) will count towards the total fours. Overthrows, all run fours and extras do not count.
- Fours scored in a super over do not count.
- In First Class games and Test Matches, only first innings fours/sixes will count.
-
Most Extras:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled in either innings due to external factors, including bad weather, unless settlement of the bet has already been determined or the match has reaches its natural conclusion without either innings being shortened to less than 80% of the originally alloted overs. Bets will stand if either innings is shortened due to it reaching a natural conclusion.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- All wide deliveries, no balls, byes, leg byes and penalty runs in the match count towards the final result. If there are runs off the bat as well as extras from the same delivery, the runs off the bat do not count towards the final total.
- Extras in a super over do not count.
- In First Class games, only first innings extras will count.
-
Most Run Outs Whilst Fielding:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, unless settlement has been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Run Outs in a super over do not count.
- In First Class games, only first innings run outs will count.
-
Most catches/keeper catches:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled in either innings due to external factors, including bad weather, unless settlement of the bet has already been determined or the match has reaches its natural conclusion without either innings being shortened to less than 80% of the originally alloted overs. Bets will stand if either innings is shortened due to it reaching a natural conclusion.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined. Catches in a super over do not count.
-
Most keeper stumpings:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled in either innings due to external factors, including bad weather, unless settlement of the bet has already been determined or the match has reaches its natural conclusion without either innings being shortened to less than 80% of the originally alloted overs. Bets will stand if either innings is shortened due to it reaching a natural conclusion.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Stumpings in a super over do not count.
-
3-way betting. Runs in over #:
- The specified over must be completed for bets to stand unless settlement has already been determined. If during the first over the innings is ended due to external factors, including bad weather, all bets will be void, unless settlement has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- In First Class matches the market refers only to each team’s first innings.
-
Highest score in First 5/6/10/12/15/20 Overs:
- If the specified number of overs are not completed in both innings the bet will be void, unless settlement of the bet has already been determined.
- In limited overs matches, if the match is shortened due to external factors after the start of the first innings, the market will be void.
- In First Class matches the market refers only to each team’s first innings.
- The market will be settles according to the official result.
-
The best result of the first partnership / Highest score at 1st dismissal:
- If the batting team reaches the end of their allotted overs, reaches their target or declares before the first wicket falls, the result will be the total amassed.
- For settlement purposes, a batsman retiring hurt does not count as a wicket.
- In limited overs matches, bets will be void if the innings has been reduced (by more than 10%)due to external factors, including bad weather, unless settlement has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- In First Class matches the market refers only to each team’s first innings.
-
Total Fours / Total Sixes / Total Boundary:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches the whole match counts. Bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Only fours/sixes/boundaries scored from the bat (off any delivery – legal or not) will count towards the total fours/sixes. Overthrows, all run fours and extras do not count.
- Fours/sixes/boundaries scored in a super over do not count.
-
Wickets Total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Retired hurt does not count as a dismissal.
-
Total runs / Team total runs:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Runs scored in super over do not count.
-
Total Ducks:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Ducks scored in super-over do not count.
-
Extras Total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- All wide deliveries, no balls, byes, leg byes and penalty runs in the match count towards the final result. If there are runs off the bat as well as extras from the same delivery, the runs off the bat do not count towards the final total.
- Extras in a super over do not count.
-
Total Run Outs:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In First Class matches the whole match counts but in the case of a drawn match bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Run outs in a super over do not count.
-
Total in highest scoring over:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- All runs, including extras, count towards settlement.
- Super overs do not count.
-
Top batter:
- In limited overs matches, bets will be void if it has not been possible to complete at least 50% of the overs scheduled to be bowled in either innings at the time the bet was placed due to external factors, including bad weather unless the innings has reached a natural conclusion.
- Top batsmen bets for First Class matches apply only to the first innings of each team, and will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined. If a player was named at the toss, but later is removed as a concussion sub, that player will still be counted, as will the replacement player. If a batsman does not bat, but was named in the starting XI, bets on that batsman will stand. If a batsman is substituted in after the in-play market has been offered, and scores the highest individual score, bets on the market will be void, unless there is a dead-heat. If a substitute (concussion, or otherwise) not named in the original XI is able to bat in their teams first innings, regardless of if they actually bat or not , bets on the market will stand.
- When two or more players score the same number of runs, these bets will settle as winners.
- Runs scored in a super over do not count.
-
Top bowler:
- In limited overs matches, bets will be void if it has not been possible to complete at least 50% of the overs scheduled to be bowled in either innings at the time the bet was placed due to external factors, including bad weather unless the innings has reached a natural conclusion.
- Top bowler bets for First Class matches apply only to the first innings of each team, and will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- If a player was named at the toss, but later is removed as a concussion sub, that player will still be counted, as will the replacement player. If a bowler does not bowl, but was named in the starting XI, bets on that bowler will stand. If a bowler is substituted in after the in-play market has been offered, and takes the most wickets, bets on the market will be void, unless there is a dead-heat. If a substitute (concussion, or otherwise) not named in the original XI bowls at least one delivery in their teams first innings,bets on the market will stand.
- If two or more bowlers have taken the same number of wickets, the bowler who has conceded the fewest runs will be the winner. If there are two or more bowlers with the same wickets taken and runs conceded, these bets will settle as winners. Wickets taken in a super over don’t count.
-
Player of the match:
- Bets will be settled on the officially declared man of the match.
- When two or more players are declared as Players of the match, these bets will settle as winners.
- If no man of the match is officially declared then all bets will be void.
-
First Ball Of Match / Second Ball Of Match / Third Ball Of Match:
- Dead balls do not count, at least one ball must be delivered or will be made void.
- Wides/No Balls/Bye/leg byes - runs allotted to these do not apply in this market e.g. 5 wides is a settled as Wide.
-
Wide balls total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to be bowled due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Wides in a super over do not count.
-
Runs total in delivery:
- The result will be determined by the number of runs added to the team total, off the specified delivery.
- For settlement purposes, all illegal balls count as deliveries. For example, if an over starts with a wide, then the first delivery will be settled as 1 and, although there has not been a legal ball bowled, the next ball will be deemed as delivery 2 for that over.
- If a delivery leads to free hit or a free hit is to be re-bowled because of an illegal delivery, the runs scored off the additional delivery do not count.
- All runs, whether off the bat or not are included. For example, a wide with three extra runs taken equates to 4 runs in total off that delivery.
-
Runs total in Over #:
- The specified over must be completed for bets to stand unless settlement has already been determined. If an innings ends during an over then that over will be deemed to be complete unless the innings is ended due to external factors, including bad weather, in which case all bets will be void, unless settlement has already been determined.
- If the over does not commence for any reason, all bets will be void.
- Extras and penalty runs will only count if they are accredited to a particular delivery within that over.
-
Wicket in Over #:
- As “Runs total in Over”.
- For settlement purposes, any wicket will count, including run outs. A batsman retiring hurt does not count as a wicket. If a batsman is timed out or retired out then the wicket is deemed to have taken place on the previous ball.
-
Over Odd/Even:
- As “Runs total in Over”.
- Zero will be deemed to be an even number.
-
Match Odd/Even:
- As “Total runs”.
-
First 5/6/10/15/20 Overs total:
- If the specified number of overs are not complete the bet will be void, unless the team is all out, declares, reaches their target or settlement of the bet has already been determined.
-
Total dismissals at first 5/10/15 Overs:
- If the specified number of overs are not complete the bet will be void, unless the team is all out, declares, reaches their target or settlement of the bet has already been determined.
- For settlement purposes, if a batsman is timed out or retired out then the wicket is deemed to have taken place on the previous ball.
-
3-way betting. Innings:
- Both teams must complete their first innings for bets to stand.
-
Innings Runs:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined. Bets placed on a future innings will remain valid regardless of the runs scored in any current or previous innings.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined. Bets will also be void in drawn first class matches, if less than 60 overs have been bowled in an incomplete innings, unless settlement of the bet has already been determined. If a team declares, that innings will be considered complete for the purposes of settlement.
- Penalty runs awarded after the conclusion of the innings, will not count towards innings runs.
-
Innings Dismissals total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Retired hurt does not count as a dismissal.
-
Innings Fours total / Innings Sixes total:
- Only fours/sixes scored from the bat (off any delivery – legal or not) will count towards the total fours/sixes. Overthrows, all run fours and extras do not count.
- Fours/sixes scored in a super over do not count.
-
Innings Extras total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined. Bets placed on a future innings will remain valid regardless of the runs scored in any current or previous innings.
- In first class or Test matches at least 60 overs must be bowled in the Innings unless team is all out or declares. All wide deliveries, no balls, byes, leg byes and penalty runs in the match count towards the final result. If there are runs off the bat as well as extras from the same delivery, the runs off the bat do not count towards the final total.
- Extras in a super over do not count.
-
Innings Run Outs total:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather, unless settlement of the bet has already been determined. Bets placed on a future innings will remain valid regardless of the runs scored in any current or previous innings.
- In first class or Test matches at least 60 overs must be bowled in the Innings unless team is all out or declares.
- Run outs in a super over do not count.
-
Innings Wide balls total:
- Wides in a super over do not count.
-
Top batter in Innings:
- The result of this market is determined on the batsman with the highest individual score in a team’s innings.
- In limited overs matches, bets will be void if it has not been possible to complete at least 50% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather.
- Top batsman bets for First Class matches apply only to the first innings of each team, and will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- If a player was named at the toss, but later is removed as a concussion sub, that player will still be counted, as will the replacement player.
- If a batsman does not bat, but was named in the starting XI, bets on that batsman will stand.
- If a batsman is substituted in after the in-play market has been offered, and scores the highest individual score, bets on the market will be void, unless there is a dead-heat.
- If a substitute (concussion, or otherwise) not named in the original XI is able to bat in their teams first innings, regardless of if they actually bat or not , bets on the market will stand.
- In the case of two or more batsmen winning with the same total the winner shall be the one who has faced the fewest deliveries, if that is the same the winner shall be who has hit the most boundaries, if there is still no clear winner then bets will be void
- Runs scored in a super over do not count.
-
Top Bowler in Innings:
- The result of this market is determined on the bowler with the highest individual number of wickets in an individual innings.
- In limited overs matches, bets will be void if it has not been possible to complete at least 50% of the overs scheduled to have been bowled at the time the bet was placed due to external factors, including bad weather.
- Top bowler bets for First Class matches apply only to the first innings of each team, and will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined. If a player was named at the toss, but later is removed as a concussion sub, that player will still be counted, as will the replacement player. If a bowler does not bowl, but was named in the starting XI, bets on that bowler will stand. If a bowler is substituted in after the in-play market has been offered, and takes the most wickets, bets on the market will be void, unless there is a dead-heat. If a substitute (concussion, or otherwise) not named in the original XI bowls at least one delivery in their teams first innings,bets on the market will stand."
- If two or more bowlers have taken the same number of wickets, the bowler who has conceded the fewest runs will be the winner. If there are two or more bowlers with the same wickets taken and runs conceded, dead heat rules will apply. Wickets taken in a super over don’t count.
- If no bowlers take a wicket in an innings then all bets will be void.
-
Total Runs Player / Player to score 25/50/75/100 runs / Highest individual score / Top batter total:
- If the batsman finishes the innings not out, as a result of a declaration, the team reaching the end of their allotted overs, or the team reaching their target; his score will be the final result. If a batsman does not bat, the bet will be void. If a batsman is not in the starting 11, bets will be void.
- If a batsman retires hurt, but returns later, the total runs scored by that batsman in the innings will count. If the batsman does not return later, the final result will be as it stood when the batsman retired.
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, unless settlement has been determined, or goes on to be determined. Result will be considered determined if the line at which the bet was placed is passed, or the batsman is dismissed.
- In drawn First Class matches, bets will be void if fewer than 200 overs are bowled, unless settlement of the bet has already been determined.
- Runs scored in a super over do not count.
-
Will score 50/100 runs:
- Same as "Total Runs Player"
-
Player Total Fours / Player Total Sixes:
- If a batsman finishes the innings not out, as a result of a declaration, the team reaching the end of their allotted overs, or the team reaching their target; his number of fours will be the final result. If a batsman does not bat, the bet will be void. If a batsman is not in the starting XI, bets will be void.
- If a batsman retires hurt, but returns later, the total fours hit by that batsman in the innings will count. If the batsman does not return later, the final result will be as it stood when the batsman retired.
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, unless settlement has been determined, or goes on to be determined. Result will be considered determined if the line at which the bet was placed is passed, or the batsman is dismissed.
- In drawn First Class matches, bets will be void if fewer than 200 overs are bowled, unless settlement of the bet has already been determined.
- Only fours/sixes scored from the bat (off any delivery – legal or not) will count towards the total fours/sixes. Overthrows, all run fours and extras do not count.
- Fours scored in a super over do not count.
-
Dismissal method (2 ways) / Dismissal method (7 ways) / Player Dismissal method:
- If the specified batsman is not out, all bets will be void.
- If the specified batsman retires, and does not return to bat later, all bets will be void. If that batsman does return to bat later and is out, bets will stand. If the specified wicket does not fall, all bets will be void.
-
Fall of # wicket:
- If the batting team reaches the end of their allotted overs, reaches their target or declares before the specified wicket falls, the result will be the total amassed.
- For settlement purposes, a batsman retiring hurt does not count as a wicket.
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, unless settlement has already been determined, or goes on to be determined. Result will be considered determined if the line at which the bet was placed is passed, or the wicket in question falls.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
-
Batters H2H:
- Bets will settle based on the official scores for the specified batsmen in the innings, as detailed in the “Total Runs Player” section above.
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, after the bet is placed unless settlement has already been determined.
-
Bowlers H2H:
- In limited overs matches, bets will be void if it has not been possible to complete at least 80% of the scheduled overs in either innings due to external factors, including bad weather, unless settlement has been determined.
- In drawn First Class matches, bets will be void if fewer than 200 overs have been bowled, unless settlement of the bet has already been determined.
- Both players must be named in the starting XI If either does not then subsequently bowl all bets are still settled.
- Wickets taken in a super over do not count.
-
Series Winner:
- Bets on Series markets stand as long as at least one scheduled match is completed, unless stated otherwise.
- If there is no further play in a Series after bets are placed, then bets will be void.
- If a Series is drawn and no draw option is offered, bets will be void.
-
Series winner after 2 Tests:
- Bets on Series markets stand as long as at least two scheduled matches is completed, unless stated otherwise.
- If there is no further play in a Series after bets are placed, then bets will be void.
-
Series Handicap:
- Bets will be void if the number of scheduled games (after bets are placed) are not played, unless settlement of bets is determined.
-
Team With Highest Innings Score:
- If a Series is drawn, bets will be void.
-
Series Correct Score:
- Series Correct Score - Bets will be void if the number of scheduled games are not played.
-
Result of the 1st Match Test/Series:
- Bets on Series markets stand as long as at least one scheduled match is completed, unless stated otherwise.
- If there is no further play in a Series after bets are placed, then bets will be void.
- If a Series is drawn and no draw option is offered, bets will be void.
-
Player Of The Series:
- Any quoted player, who takes no part in the specified series, will be void. Winner as declared by the ICC. Dead-heat rules apply.
-
Double chance:
- All match betting will be settled in accordance with official results.
- If there is no official result, all bets will be void.
-
Players sent off / retired out:
- A player being sent off is viewed as retired out, so will be settled as a wicket.
-
Concussion substitutions:
- When a player leaves the field as a concussion substitute, this will not count as a wicket. If the player does not return later, the final result will be as it stood when the player left the field. When a player enters the match as a concussion substitute, for settlement purposes both they and the player replaced will be looked upon as to have played a full part in the match.
-
Penalty runs after the conclusion of an innings:
- Penalty runs added to a team’s total after the start of the other team’s innings will not count towards settlement of markets in the previous innings
Simulated reality league
- Simulator of real sports tournaments (time, format of matches (football: 2 halves of 45 minutes, hockey: 3 periods of 20 minutes, etc.), the rules of the competition)
- In case of technical the graphic broadcast (pitch animation) of the event (disconnection, DDoS, etc.) is interrupted, this is not a reason for canceling bets, except in situations in which it is impossible find out the final result of the match.
-
Official source for bets settlement
Virtual football
- Virtual Football League (VFL) is composed of 16 teams and continuously occurring seasons. Each season consists of 30 virtual matchdays (home and away matches).
-
Features:
- A season lasts 141 minutes and includes 30 tours;
- A tour consists of eight matches;
- The duration of a match is 4 minutes 35 seconds;
- Each match is composed of: "Pre Match" period; "First half"; "Second half"; "Post Match" period.
- "Pre Match" period heralds the beginning of the match and lasts 60 seconds. The duration of each half is 45 minutes, as in real football. But only pieces of dangerous moments with a total duration of 1 minute 30 seconds are shown. Break between halves is 10 seconds. At the end of each meeting, 10 seconds are given to "Post Match" period, and 15 seconds to complete the "Virtual Tour of the game."
- Acceptance of bets on VFL matches stops in 10 seconds before the start of the match. Bets can be placed on all upcoming matches of the current season. Each "matchday" shows all selected matches of the tour at the current odds.
-
Types of wagers available for betting:
- Regular time (Win Team 1; Draw, Win Team 2);
- Handicap (H1, H2);
- The first half (Team 1 Win, Draw, Win Team 2);
- Total (over; under);
- Correct score (from 0:0 to 3:3; other);
- Who will score the next goal at the score 0:0? (Team 1; None; Team 2).
- The bets are settled immediately after the match.
- The result of each event is not determined by us, neither it depends on us, bettors or placed bets.
- If the event did not take place due to failure of the application and in the case of such errors as incorrect displaying of wagers and odds - the bets will be refunded.
- Connection failure, communication disruptions, browser errors, windows closing and other technical reasons are not grounds for bet cancellation as they do not affect the course of a match, result and the settlement of bets. In addition, delays in the broadcasting are not the reason to cancel bets.
- The bettor is responsible for proper selection of bets.
Virtual Basketball
- Virtual Basketball League (VBL) is composed of 16 teams and continuously occurring seasons. Each season consists of 30 virtual match days (home and away games).
-
Features:
- A season lasts 108 minutes and includes 30 rounds;
- A round consists of eight matches;
- The duration of a match is 3 minutes 30 seconds;
- Each match is composed of: "Pre Match" period; "Match"; "Post Match" period.
- "Pre Match" period heralds the beginning of the match and lasts 15 seconds. The duration of each match is 2 minutes 40 seconds. Only pieces of dangerous moments with a total duration of 1 minute 30 seconds are shown. At the end of each match, 35 seconds are given to "Post Match" period.
- Acceptance of bets on VBL matches stops in 10 seconds before the start of the match. Bets can be placed on all upcoming matches of the current season. Each "match day" shows all selected matches of the tour at the current odds.
-
Types of wagers available for betting:
- Fulltime result (Win Team 1; Draw, Win Team 2);
- Handicap (H1, H2);
- The first half (Team 1 Win, Draw, Win Team 2, Handicap H1; H2, Total over; under);
- Total (over; under);
- Individual total.
- The bets are settled immediately after the match.
- The result of each event is not determined by us, neither it depends on us, bettors or placed bets.
- If the event did not take place due to failure of the application and in the case of such errors as incorrect displaying of bets and odds - the bets will be refunded.
- Connection failure, communication disruptions, browser errors, windows closing and other technical reasons are not grounds for bet cancellation as they do not affect the course of a match, result and the settlement of bets. In addition, delays in the broadcasting are not the reason to cancel bets.
- The bettor is responsible for proper selection of bets.
Virtual Tennis
- Virtual tennis tournament is composed of two Grandstand Open and Britannia Open, each including participation of 16 players.
-
Features:
- one tournament consists of 15 matches and goes in 4 rounds: 1/8 finals, 1/4 finals, semifinals, finals;
- duration of one tournament is 26 minutes, rounds of each tournament change one by one:
-
1/8 Britannia Open
Grandstand Open final
1/4 Britannia Open
1/8 Grandstand Open
Britannia Open semifinals
1/4 Grandstand Open
Britannia Open final
Grandstand Open semifinals
- duration of one round / match is 3 min 45 sec;
- matches are played with 3 sets format (until victory in two sets).
- The bets for the upcoming Grandstand Open round are accepted during the appropriate Britannia Open tour and vice versa.
-
Types of wager available for placing bets:
- Fulltime result (W1 - winning of the 1st player, W2 - winning of the 2nd player);
- Total (over, under) - settled by games;
- Exact score (2:0, 2:1, 0:2, 1:2);
- Winning the 1st set (W1 - winning of the 1st player, W2 - winning of the 2nd player).
- 1st set total (over, under) - settled by games in the first set.
- Bets are settled immediately after the match.
- The result of each event is not determined by us, neither it depends on us, bettors or placed bet amount.
- If the event did not take place due to failure of the application and in the case of such errors as incorrect displaying of bets and odds - the bets will be refunded.
- Connection failure, communication disruptions, browser errors, windows closing and other technical reasons are not grounds for bet cancellation as they do not affect the course of a match, result and the settlement of bets. In addition, delays in the broadcasting are not the reason to cancel bets.
- The bettor is responsible for proper selection of bets.
Wager for the deposit
We can apply restrictions on the withdrawal of money from your account by setting a wager on the deposit. This means that it is possible to withdraw money from the account only if the user spent part of the deposited money on bets. The size of the Wager may vary. Information about the wager is available in the account section.
Example 1. A wager with a coefficient of 0.5 is set for a user.
Wager condition: 50% of the deposit must be spent on bets.
- user deposits $10
- wager installed: $5
- user made a bet on $3
- user wins $5
Withdrawing money is not possible, because it is necessary to place additional bets for $2
Example 2. A wager with a coefficient of 0.5 is set for a user.
Wager condition: 50% of the deposit must be spent on bets.
- user deposits $10
- wager installed: $5
- user made a bet on $5
- user wins $15
Withdrawal of money is possible for an amount not exceeding the state of the balance, i.e. $20 ($10 deposit - $5 bet + $15 win)


